OpenStudy (dancing_stars23):

1.The relationship between pressure and temperature of a gas, when volume and moles of a gas are held constant, is: P*T = k. True False 2. The relationship between pressure and volume, when moles and temperature of a gas are held constant, is: PV = k. We could say then, that: If the volume of the gas is doubled, the pressure also doubles. If the volume of a gas is halved, the pressure is doubled. If the volume of gas is doubled, the pressure is unaffected. If the volume of a gas is halved, the pressure is tripled.

OpenStudy (matt101):

For the first question, think of the gas law: PV=nRT. If you rearrange the equation so that P and T are on the same side, you get P/T = nR/V. If moles and volume are constant (R is already a constant, then P/T = k. So what's the answer for #1? For the second question, you can do a quick check to see what is right. If the first option is right, then 2P x 2V should equal PV (because the product is supposed to be constant). But 2P x 2V = 4PV...four times what it should be! So that can't be right. Work through the other options to see which answer is correct!

OpenStudy (dancing_stars23):

@matt101 true? and number#2?

OpenStudy (matt101):

Is P*T the same as saying P/T?

OpenStudy (dancing_stars23):

nooooo @matt101

OpenStudy (thadyoung):

Yes, they are proportional. Think about the ideal gas law formula: \(\sf \color{red}{PV = nRT}\) pressure is proportional to temperature.

OpenStudy (matt101):

So @dancing_stars23 is the answer true or false?

OpenStudy (dancing_stars23):

@matt101 the answer would have to be false then

OpenStudy (matt101):

Right! Now for the second question, reread my answer above and tell me which answer you think is right.